A Hedge fund in 17 financed their $100 million of MBS by using seven-day repurchase agreements in addition to their investors’ capital. They borrow the maximum amount, haircut is 10%, and the interest rate is 2% per year, which of the following is closest to how much interest they will owe at the end of the first seven-day term? (a) $30,000 (b) $35,000 (c) $38,000 (d) $40,000 (e) None of these are within $5,000
Answer #1

Answer is (b) $35,000.

Maximum borrowing amount = value financed*(1 - haircut rate) = $100 million*(1 - 0.10) = $100 million*0.90 = $90 million

interest owed = Maximum borrowing amount*annual interest rate*repo term/360 days = $90,000,000*2%*7/360 = $1,800,000*7/360 = $12,600,000/360 = $35,000

they will owe $35,000 interest at the end of the first seven-day term.